ON ARITHMETIC PROGRESSIONS ON ELLIPTIC CURVES

ANDREW BREMNER

Abstract. We study arithmetic progressions in the x-coordinates of rational points on elliptic curves. An infinite family of elliptic curves is found, each containing an arithmetic progression of length 8.